RC Update and Run Plan for July 30, 2020
- 1 RC Meeting Information
- 2 General updates
- 3 Current Run Plan
- 4 Notices in Effect for Shift Crew
- 5 Planned Accelerator Configuration Changes
- 6 Useful Links
RC Meeting Information
RC Daily Meetings are at 4:15 PM, 2nd floor Counting House Meeting Room
Thursday July 30 2020.
- Hall C sweep at 3 pm; after that Hall C in Controlled access (at some point in Power Permit for PSS re-cert: all people must be out of the hall for that)
- The Hall is ready; downstream turbo pump has been replaced - so far so good. I will check on the pump tomorrow morning to make sure it didn't trip.
- Projection from this morning for beam arrival into Hall C: Friday day shift
- Accelerator: BSY3C arc power supply (if this doesn't get fixed, no beam to Hall C) https://logbooks.jlab.org/entry/3818113
Recent polarization measurements
- Longitudinal (180): 52% (I am citing Junhao)
Current Run Plan
First beam into the Hall
- Tune beam is allowed; make sure the target ladder is in "no target" position - MCC should notify the RC one hour prior to tune beam arrival in Hall C
- Before allowing any CW into Hall C we have to go through the beam centering procedure to establish optimal beam positions for the 3He cell and for the Reference cell to make sure there is no beam strike in the target region - expert driven (Hall C RC will call Brad when we are at that step)
- After the optimal beam positions are established MCC will take over the program to certify the machine tune; the target ladder must be in the "no target position" and the position lock must be engaged. The TO/SL must monitor that the beam positions stay at nominal values. We expect 30-40 muA to the dump as part of the "machine tune certification" exercise
- Once MCC will be ready to send Physics beam follow the plan outlined here:
Normal Production running
- If beam goes away for more than 2h, or if during the downtime, MCC does operations that may change the beam tuning (except work on transport magnets), the do beam recovery procedure (do YAG viewer part only if your are suspicious of the harp scan results)
- Make sure all the beam specs are followed (e.g. no raster size below 4mm on 3He cell).
- Check beam parameters/configuration and post an entry like this once a shift
- Always take 120Hz of EDTM
- Fill the BTA with realistic numbers.
- General idea: take x hours of data....
Notes for tomorrow's 7:45 meeting
- Ask Tony dela Cruz about tuning BCM1,2
- Reiterate about run conditions on cell targets: raster always on, ramp rate 1muA/sec, beam position lock on
When there is a down:
- If the down is 2+ hours call Simona: she will check with the RC to see if there is an opportunity for work in the Hall (SoLID Cherenkov parasitic test)
Notices in Effect for Shift Crew
- Policy for CANS-related hall status change: Make a logbook entry when the hall goes in, or comes out of laser-controlled access.
- Magnet policies
- Spectrometer GUI (
go_magnets) should always be visible on Shift Leader computer, hcdesk3
- HMS Q1 must be ramped down in 50A steps in all cases
- Do not ramp HMS Q2 to zero, only go to 10A and then switch off power.
- If any of the Spectrometers magnets trip, take a screenshot of the PSU window before resetting the interlock and post it to hclog.
- Spectrometer GUI (
- Inform the RC of any access that the Target Team requires. The RC will coordinate with RadCon to make everything run smoothly.
- Please make sure that the orbit locks are enabled and the beam does not drift from its nominal position.
- Ramp rate lock instructions are here and added to checklist
Planned Accelerator Configuration Changes
- Bluejeans remote connection information:
To join the Meeting: https://bluejeans.com/861439920 To join via Room System: Video Conferencing System: bjn.vc -or-126.96.36.199 Meeting ID : 861439920 To join via phone : 1) Dial: +1.408.740.7256 (US (San Jose)) +1.888.240.2560 (US Toll Free)) +1.408.317.9253 (US (Primary, San Jose)) (see all numbers - http://bluejeans.com/numbers) 2) Enter Conference ID : 861439920